# Comparing matching prescriptions between pre-equilibrium and hydrodynamic models in high-energy nuclear collisions

### v1

@@ II.2 Matching the pre-equilibrium model to fluid dynamics
On the one hand, we use the by now standard recipe, consistent with Landau matching, in which the energy density $e_{\rm K}(\mathbf{x})$ at a given transverse position (in practice, in a cell) at the end of the pre-equilibrium model, here KøMPøST, becomes the energy density $e_{\rm M}(\mathbf{x})$ at the same point in the hydrodynamic code, namely MUSIC: $e_{\rm M}(\mathbf{x})=e_{\rm K}(\mathbf{x})$. In the following, this approach will be referred to as “energy matching”. Since we also feed the final velocity of KøMPøST as is into MUSIC, this mode ensures local energy conservation. Eventually, the bulk viscosity in MUSIC is initialized as
@@ III Results
Figure 1: Time dependence of the energy (left) and (equivalent) equilibrium entropy (right) per unit rapidity, using the energy-matching (blue) and entropy-matching (green) prescriptions. The vertical dotted line stands for the switching time $\tau_{\rm hydro}$ between KøMPøST and MUSIC. The top plots are for the average central (0–10%) event and the bottom ones are for the average event in the 40–50% centrality class.
@@ III.1 Thermodynamic and fluid-dynamical quantities around the switching time
Figure 2: Time dependence of the thermodynamic pressure (left) and bulk pressure (right) per unit rapidity, using the energy-matching (blue) and entropy-matching (green) prescriptions, for the average central (0–10%, top) and peripheral (40–50%, bottom) events. The vertical dotted line stands for the switching time $\tau_{\rm hydro}$ between KøMPøST and MUSIC.
@@
In contrast, the mismatch between the equations of state in KøMPøST and MUSIC means that there are discontinuities at the switching time. Thus, energy is discontinuous when applying entropy matching (left, green curves), while reciprocally the integral of $s_{\rm eq.}$ is discontinuous in the simulation with energy matching (right, blue curves). Due to the different initial $K$-factors, the values of the energy — and thus the equivalent equilibrium entropy — necessarily differ in the two matching prescriptions in the pre-equilibrium stage. Yet for those quantities with a jump at $\tau_{\rm hydro}$, the values may become closer in MUSIC, because the final charged multiplicity should be the same. This trend is not so clear in the peripheral events, but one should pay attention that the difference between the values in MUSIC after using the two prescriptions are actually of the same absolute size (about 30 GeV for energy) as in central events. Furthermore, one should keep in mind that the integral of the equilibrium entropy $s_{\rm eq.}$ is not the total entropy (per unit rapidity) in the (MUSIC) fluid. In addition to $s_{\rm eq.}$, one should also consider the non-equilibrium contributions from shear stresses and bulk pressure [33, 34] — and as we shall discuss next, bulk pressure in MUSIC does depend on the matching recipe. This also explains why the slight decrease of the integral of $s_{\rm eq.}$ in the MUSIC stage is not in contradiction with the second law of thermodynamics. In fact, there are also parts of the system that have already particlized, and which we do not take into account in those plots.
In Fig. 2, we show the evolution of thermodynamic (left) and bulk pressure (right) integrated over the transverse plane. We recall that in KøMPøST $P=e/3$ and $\Pi=0$, so that the portions of these plots before $\tau_{\rm hydro}$ are either degenerate with Fig. 1 or trivial. Indeed, the most interesting is the behavior at $\tau_{\rm hydro}$, where these pressures are discontinuous. In all cases, the jumps in $P$ or $\Pi$ are smaller with the entropy-matching prescription than with energy matching. When equating $s_{\rm eq.}$ in MUSIC and KøMPøST, the energy density at the start of MUSIC is larger than at the end of KøMPøST: this corresponds to a higher temperature, so that the LQCD equation of state is closer to being conformal, as it is in KøMPøST, which explains why the gap in thermodynamic pressure is smaller. In turn, since the sum $P+\Pi$ is continuous this results of an equally smaller gap in bulk pressure, such that the initial $\Pi$ in MUSIC is smaller with entropy matching.
@@ III.2 Final-state observables
As already mentioned, the final charged-hadron multiplicity was fixed to determine the initial normalization for each approach, so that $N_{\rm ch.}$ is by construction the same with both matching prescriptions. We also computed a number of other observables in the final state of the SMASH evolution, [^5] For that analysis, we used the SPARKX package [39]. to test how much they are affected by the change in matching recipe. More precisely, we looked at the average over events of the first five anisotropic-flow harmonics $v_{n}$ and of the average transverse momentum $[p_{T}]$ of charged hadrons. We also looked at an observable that is more sensitive to event-by-event fluctuation, namely the linear Pearson coefficient $\rho([p_{t}],v_{2}^{2})$ for the correlation between $[p_{T}]$ and $v_{2}^{2}$ [40]. For the flow coefficients we computed the modulus of the average of ${\rm e}^{{\rm i}n\varphi}$ over the charged hadrons in an event, where $\varphi$ is the particle azimuthal angle with respect to the known impact-parameter direction.
Table 1: Final-state observables for 5000 events in the centrality class 0–10%.
@@
The values of these observables for central collisions are shown in Tab. 1 — the trend in peripheral collisions is the same and not shown. One sees that the values of the observables are very similar with both matching prescriptions, mostly at the percent level. The only exceptions are $v_{1}$, which varies by about 3%, and especially $\rho([p_{t}],v_{2}^{2})$, which changes by almost 10%.
From these few results, we would conclude that the precise matching prescription used has little impact on “global” observables averaged over events. More differential observables, in particular those sensitive to event-by-event fluctuations, may be more impacted. However, for the 5000 events in the 40–50% centrality class the Pearson coefficient $\rho([p_{t}],v_{2}^{2})$ changes from 0.1708 with energy matching to 0.1773 with entropy matching, i.e. by only a few percent. Note that our using the regulators in MUSIC may partly suppress the influence of the different bulk pressures (Fig. 2, right panel) in the hydrodynamic stage on final-state observables.
@@ IV Discussion
It is clear that the proposed matching recipe has the unpleasant feature that it does not conserve energy at the interface between pre-equilibrium and fluid dynamics. This is admittedly inelegant, but one should remember that in the present state of the art, the absolute energy density at the beginning of fluid dynamics is determined from the multiplicity in the final state of the evolution, from which an ad hoc $K$-factor in the initial state (or sometimes at the beginning of hydrodynamics [15]) is deduced. Since multiplicity is rather associated with total entropy, a prescription based on the latter would seem more motivated [^6] See Ref. [41] for a discussion of whether energy or entropy density should be used in the initial state of fluid dynamics. — at least as long there is no initial-state model that predicts the value of the initial energy without any indeterminate $K$-factor. Now, our approach in the present study does not exactly conserve entropy at the switch between KøMPøST and MUSIC, but it is already a significant step in that direction, since the jump in total entropy is only of second order in the gradients, as discussed at the end of Sec. II.

### v2

@@ II.2 Matching the pre-equilibrium model to fluid dynamics
On the one hand, we use the by now standard recipe, consistent with Landau matching, in which the energy density $e_{\rm K}(\mathbf{x})$ at a given transverse position (in practice, in a cell) at the end of the pre-equilibrium model, here KøMPøST, becomes the energy density $e_{\rm M}(\mathbf{x})$ at the same point in the hydrodynamic code, namely MUSIC: $e_{\rm M}(\mathbf{x})=e_{\rm K}(\mathbf{x})$. In the following, this approach will be referred to as “energy matching”. Since we also feed the final velocity of KøMPøST as is into MUSIC, this mode ensures local energy conservation. Eventually, the bulk pressure in MUSIC is initialized as
@@ III Results
Figure 1: Time dependence of the energy (left) and (equivalent) equilibrium entropy (right) per unit rapidity, using the energy-matching (blue full line) and entropy-matching (red dashed line) prescriptions. The vertical dotted line stands for the switching time $\tau_{\rm hydro}$ between KøMPøST and MUSIC. The top plots are for the average central (0–10%) event and the bottom ones are for the average event in the 40–50% centrality class.
@@ III.1 Thermodynamic and fluid-dynamical quantities around the switching time
Figure 2: Time dependence of the thermodynamic pressure (left) and bulk pressure (right) per unit rapidity, using the energy-matching (blue, full) and entropy-matching (red, dashed) prescriptions, for the average central (0–10%, top) and peripheral (40–50%, bottom) events. The vertical dotted line stands for the switching time $\tau_{\rm hydro}$ between KøMPøST and MUSIC.
@@
In contrast, the mismatch between the equations of state in KøMPøST and MUSIC means that there are discontinuities at the switching time. Thus, energy is discontinuous when applying entropy matching (left, red curves), while reciprocally the integral of $s_{\rm eq.}$ is discontinuous in the simulation with energy matching (right, blue curves). Due to the different initial $K$-factors, the values of the energy — and thus the equivalent equilibrium entropy — necessarily differ in the two matching prescriptions in the pre-equilibrium stage. Yet for those quantities with a jump at $\tau_{\rm hydro}$, the values may become closer in MUSIC, because the final charged multiplicity should be the same. This trend is not so clear in the peripheral events, but one should pay attention that the difference between the values in MUSIC after using the two prescriptions are actually of the same absolute size (about 30 GeV for energy) as in central events. Furthermore, one should keep in mind that the integral of the equilibrium entropy $s_{\rm eq.}$ is not the total entropy (per unit rapidity) in the (MUSIC) fluid. In addition to $s_{\rm eq.}$, one should also consider the non-equilibrium contributions from shear stresses and bulk pressure [33, 34] — and as we shall discuss next, bulk pressure in MUSIC does depend on the matching recipe. This also explains why the slight decrease of the integral of $s_{\rm eq.}$ in the MUSIC stage is not in contradiction with the second law of thermodynamics. In fact, there are also parts of the system that have already particlized, and which we do not take into account in those plots.
In Fig. 2, we show the evolution of thermodynamic (left) and bulk pressure (right) integrated over the transverse plane. We recall that in KøMPøST $P=e/3$ and $\Pi=0$, so that the portions of these plots before $\tau_{\rm hydro}$ are either degenerate with Fig. 1 or trivial. Indeed, the most interesting is the behavior at $\tau_{\rm hydro}$, where these pressures are discontinuous. In all cases, the jumps in $P$ or $\Pi$ are smaller with the entropy-matching prescription than with energy matching. When equating $s_{\rm eq.}$ in MUSIC and KøMPøST, the energy density at the start of MUSIC is larger than at the end of KøMPøST: this corresponds to a higher temperature, so that the LQCD equation of state is closer to being conformal, as it is in KøMPøST, which explains why the gap in thermodynamic pressure is smaller. In turn, since the sum $P+\Pi$ is continuous this results in an equally smaller gap in bulk pressure, such that the initial $\Pi$ in MUSIC is smaller with entropy matching.
@@ III.2 Final-state observables
As already mentioned, the final charged-hadron multiplicity was fixed to determine the initial normalization for each approach, so that $N_{\rm ch.}$ is by construction almost the same with both matching prescriptions. We also computed a number of other observables in the final state of the SMASH evolution, [^5] For that analysis, we used the SPARKX package [39]. to test how much they are affected by the change in matching recipe. The reported observables are integrated over all particles at mid-(pseudo)rapidity, $|\eta|<0.5$, with no cut on transverse momentum. More precisely, we looked at the average over events of the first five anisotropic-flow harmonics $v_{n}$ and of the average transverse momentum $[p_{T}]$ of charged hadrons. We also looked at an observable that is more sensitive to event-by-event fluctuations, namely the linear Pearson coefficient $\rho([p_{T}],v_{2}^{2})$ for the correlation between $[p_{T}]$ and $v_{2}^{2}$ [40]. For the flow coefficients we computed the modulus of the average of ${\rm e}^{{\rm i}n\varphi}$ over the charged hadrons in an event, where $\varphi$ is the particle azimuthal angle with respect to the known impact-parameter direction.
Table 1: Final-state observables at mid-rapidity for 5000 events in the centrality class 0–10%.
@@
The values of these observables for central collisions are shown in Tab. 1 for central events. One sees that the values of the observables are very similar with both matching prescriptions, mostly at the percent level. The only exceptions are $v_{1}$, which varies by about 3%, and especially $\rho([p_{T}],v_{2}^{2})$, which changes by almost 10%.
Table 2: Final-state observables at mid-rapidity for 5000 events in the centrality class 40–50%.
Observable Energy matching Entropy matching
$\expectationvalue{N_{\rm ch.}}$ $259$ $282$
$\expectationvalue{v_{1}}$ (%) 1.476 1.399
$\expectationvalue{v_{2}}(\%)$ 7.247 7.309
$\expectationvalue{v_{3}}$ (%) 2.578 2.621
$\expectationvalue{v_{4}}$ (%) 1.563 1.553
$\expectationvalue{v_{5}}$ (%) 1.327 1.274
$\expectationvalue{[p_{T}]}$ (MeV) 631.21 629.21
$\rho([p_{T}],v_{2}^{2})$ 0.1708 0.1773
From these few results, we would conclude that the precise matching prescription used has little impact on “global” observables averaged over events. More differential observables, in particular those sensitive to event-by-event fluctuations, may be more impacted. However, for the 5000 events in the 40–50% centrality class (see Tab. 2) the Pearson coefficient $\rho([p_{T}],v_{2}^{2})$ changes from 0.1708 with energy matching to 0.1773 with entropy matching, i.e. by only a few percent. In contrast, the average charged multiplicity now changes by almost 10 percent, although both values remain in good agreement with the ALICE data. Note that our using the regulators in MUSIC may partly suppress the influence of the different bulk pressures (Fig. 2, right panel) in the hydrodynamic stage on final-state observables.
@@ IV Discussion
It is clear that the proposed matching recipe has the unpleasant feature that it does not conserve energy at the interface between pre-equilibrium and fluid dynamics. This is admittedly inelegant, but one should remember that in the present state of the art, the absolute energy density at the beginning of fluid dynamics is determined from the multiplicity in the final state of the evolution, from which an ad hoc $K$-factor in the initial state (or sometimes at the beginning of hydrodynamics [15]) is deduced. Since multiplicity is rather associated with total entropy, a prescription based on the latter would seem more motivated [^6] See Ref. [41] for a discussion of whether energy or entropy density should be used in the initial state of fluid dynamics. — at least as long as there is no initial-state model that predicts the value of the initial energy without any indeterminate $K$-factor. Now, our approach in the present study does not exactly conserve entropy at the switch between KøMPøST and MUSIC, but it is already a significant step in that direction, since the jump in total entropy is only of second order in the gradients, as discussed at the end of Sec. II.